OpenStudy (anonymous):

Check My answers!! Algebra 1. 4-t=3(t-1)-5 a. 3 b. 6 c. no solution (my answer) d. identity 2. 8x-2(x+1)=2(3x-1) a. 0 b. 2 c. no solution d. identity(my answer) 3. 3(c-2)=2(c-6) a. 2 b. -6(my answer) c. no solution d. identity 4. 0.5(m+4)=3(m-1) a.-2 b. 2(my answer) c. no solution d. identity Thanks!

OpenStudy (anonymous):

first one is wrong you can solve \[4-t=3(t-1)-5\]

OpenStudy (anonymous):

second one is right \[ 8x-2(x+1)=2(3x-1)\] both sides are equal to \[6x-2\]

OpenStudy (anonymous):

third one is right

OpenStudy (anonymous):

fourth one is right only the first is wrong

OpenStudy (anonymous):

you will only get "no solution" if the variable drops out and you are left with something like \(3=5\) \[ 4-t=3(t-1)-5\]multiply out on the left \[4-t=3t-3-5\]combine like terms \[4-t=3t-8\]add \(t\) \[4=4t-8\]add \(8\) \[12=4t\]divide by \(4\) \[3=t\]

OpenStudy (anonymous):

i meant multiply out on the RIGHT i suppose

OpenStudy (anonymous):

Thanks I get it now!

OpenStudy (anonymous):

yw